OpenStudy (anonymous):

The height of a trapezoid is 13 inches, one of the bases measures 22 inches. If the area of the trapezoid is 318.5 square inches, find the length of the other base. A. 2.5 inches B. 71 inches C. 34.25 inches D. 27 inches

OpenStudy (***[isuru]***):

\[Area = \frac{ A + B }{ 2} \times height\] 

OpenStudy (***[isuru]***):

so... \[318.5 = \frac{ B + 22 }{ 2 } \times 13\] solve it for B can u do it ?

OpenStudy (anonymous):

159.25bx^13+3503.5x^13

OpenStudy (***[isuru]***):

( B + 22) x13 = 318.5 x2 = 637 (B + 22) = 637/13 = 49 (B + 22) = 49 B = 27 got it :) ?